Systems Staff Meeting Notes
November 30, 2007
November 30, 2007
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Chris Nicolich, Amy Polk, Stephen WilliamsAnnouncements
Review Minutes of 11/16/07
- Stephen will be painting his new condo over the weekend.
- Effort Tracking needs to be returned to HR.
- Guitar Hero to be played in Library West on December 13 as a break from finals for students.
Ongoing Systems Updates
- Watercooler listserv trouble ticket to be reassigned to Will.
- Orders - hope to get caught up today.
- DLC - switch and network card to be ordered.
- We will be turning off the maintenance mode for Library West for two weeks due to extended hours (finals).
Software Projects Status
- Will is working on the new EPolicy for McAfee. The new version of McAfee is working better - no Trouble Tickets reporting problems as of late.
- SAN - server is configured. Virtual server has been installed. The management server database connects to NT18. It has two arrays: one for DLC, one for general Library use. The DLC array is configured for multimedia. Logan will be working on the switches today - then storage space. After that, he will move on to MSL.
- Additional items to be ordered for the Maps project - need to get a quote.
- SMS - Amy would like to go to training for it. She may need to look into using the newer version of it.
- 3480 tape conversion - all cartridges have been converted to VTS. Need to revisit a few problem tapes and compare the non-labeled tapes that were converted.
- Backup Exec upgrade - have experienced much difficulty with this upgrade. As it turns out, continuous protection for Exchange is not possible with clustered servers.
- Smathers2K3SQL was rebooted this morning. No errors - had to delete a Backup Exec remote agent folder.
- Logan is working on consolidated logging for Linux boxes. Will keep for 30-60 days. Security logs only. Runs on the VM server.
- Problems finding Smatherslib - Charlotte Barnes had the problem earlier this week. Will removed/added to the domain and changed the DNS settings - did not fix. He added the WINS server back and that was the fix. Need to make sure WINS server is listed in the master browser.
Systems Security Management
- Blog/wiki meeting with Marilyn this week - she wanted to use a wiki to teach a class this spring. She was discouraged from a wiki as it would be giving too broad of access and we suggested she use a commercial option. She asked if there is a Library policy on the matter. Blog server is up and running - waiting for a template from Tom. Will need to contact Marcus Morgan at CNS to make it live. Policy needs to be set. Will approach the Web 2.0 committee on this topic.
- Grover - working through the ticket page and security. Resolving policies in combining tickets and SysRequest.
- UFDC - Mark is working on other tools.
- Tracking database - Tom is working on importer and database redesign.
- Stephen Williams worked with Lori on problems with the carrel reservation system.
- Lori sent an email asking for an analysis of checkout data for the last several years. Files need to be sent to OCLC. Huge file transfers.
- DLC - Mark wanted to upgrade the Greenstone boxes. Logan upgraded the Development server, but is getting some errors. He is working on it. It is running on the new version without encryption. SCSI card for the Development server has been received. Will install soon. They want the monitoring tool which will warn when they are reaching capacity.
- Meeting held.